OpenStudy (anonymous):

y'- 3y = x.e^(2x) + 6 with undetermined coef method any help ???

OpenStudy (anonymous):

first solve the homogenious equation y'-3y=0

OpenStudy (anonymous):

this can be put in as m-3=0

OpenStudy (anonymous):

m=3 k=1

OpenStudy (anonymous):

\[y_c=c_1e^{3x}\]

OpenStudy (anonymous):

Yeah so true

OpenStudy (anonymous):

@mukushla .. can you finish this. I gotta go do something

OpenStudy (turingtest):

@mukushla is says to do this by undetermined coefficients I think the correct guess for the particular is \[Y_p=(Ax+B)e^{2x}+C\]

OpenStudy (anonymous):

@mukushla if u do not want to do it , do not do it ok?

OpenStudy (anonymous):

@TuringTest you are right but i could not understand too much, can u explain more please ?

OpenStudy (anonymous):

@SkykhanFalcon ok

OpenStudy (turingtest):

plug that guess into the original equation

OpenStudy (turingtest):

\[Y_p=(Ax+B)e^{2x}+C\]\[Y_p'=2(Ax+B)e^{2x}+Ae^{2x}\]\[Y_p''=4(Ax+b)e^{2x}+4Ae^{2x}\]

OpenStudy (anonymous):

that's correct turing

OpenStudy (turingtest):

\[Y_p''-3Y_p=xe^{2x}+6\]\[4Axe^{2x}+4Be^{2x}+4Ae^{2x}-3Axe^{2x}-3Be^{2x}-3C=xe^{2x}+6\]thanks for double-checking :)

OpenStudy (anonymous):

@TuringTest i love u dude :D thank u so much

OpenStudy (anonymous):

so true and so helpful

OpenStudy (turingtest):

\[Axe^{2x}+Be^{2x}-3C=xe^{2x}+6\]from which we can infer\[A=1\]\[B=0\]\[-3C=6\implies C=-2\]so\[Y_p=xe^{2x}-2\]I hope....

OpenStudy (anonymous):

@TuringTest can i ask a question

OpenStudy (anonymous):

you do not understand the \[(Ax+b)e^x+c\]? your particular solution will be in the form of your f(x) on the right side since you have xe^x, you have a polynomial of power one and e^x therefore the base for that would be \[(Ax+B)e^x\] lastly you have 6 which is a c so your particular will be in the form \[(Ax+B)e^x+c\] this is your particular solution wwith undetermined coefficients. since it's a solution , you can sub it into the equation like turing did and solve as such

OpenStudy (turingtest):

oh crap it's y' not y''

OpenStudy (turingtest):

I messed up by going to far :S

OpenStudy (anonymous):

lol o well haha

OpenStudy (anonymous):

@TuringTest how can you know Yp= (Ax+B)e^2x +C ???

OpenStudy (anonymous):

I just explained it Skykhan however i can give you a table that shows some of the ways...

OpenStudy (anonymous):

oo Thank u

OpenStudy (turingtest):

\[Y_p'-3Y_p=xe^{2x}\]\[2(Ax+B)e^{2x}+Ae^{2x}-3Axe^{2x}-3Be^{2x}-3C=xe^{2x}+6\]\[-Axe^{2x}+(A-B)e^{2x}-3C=xe^{2x}+6\]hope I did it right that time